# Phase space compression of a positive muon beam in two spatial dimensions

### v1

@@ 4 Experimental setup
Figure 4 (Top) shows the placement of the various scintillators around the target as side (Left) and top (Right) views. These scintillators (positron detectors) can be classified into two main categories: the transverse detectors (Trans 2 and Trans 3) used to expose the muons’ motion along the -axis, and the Tile detectors that are sensitive to the muons reaching the target tip exposing at the same time longitudinal () and transverse () compressions. For reference the Trans detectors have a size of () mm3 while the bottom Tile detectors are () mm3.
Figure 4 (Middle) and (Bottom) depict the detection efficiencies of these detectors, simulated with GEANT4, projected onto the and -plane, respectively. The position-dependence of the detection probability, which results in a relatively good position resolution, was obtained by embedding the scintillators into a massive copper collimator and by organizing the scintillators in telescope-pairs read out in coincidence (top and bottom detectors). Each detector effectively covers distinct zones of the target volume. For instance, the Tile detectors are only sensitive to muon-decays occurring at the tip of the target at their respective -position. For reference Fig. 4 also indicates the position of the muons at time , i.e., the muon stopping distribution before compression takes place. To further increase the position resolution an energy threshold is applied to each positron counter, efficiently removing events from secondary particles. Each scintillator was optically coupled to a wavelength-shifting fiber to transport their optical signals from the cryogenic section of the setup to outside the cryostat. The fibers were then read out by SiPMs [28].
@@ 5 Measured time spectra and comparison to simulation
The time spectra shown in Fig. 5 clearly prove that starting from the muon stopping distribution centered at mm (see Fig. 4) at , the muons gradually drift in the -direction over time. At , the muons are located in the muon stopping region, which falls outside the detection range of the various detectors, resulting in zero or nearly zero counts in all three time spectra. Around µs, the muons pass through the region with the maximum acceptance for the Trans 3 detector, as evidenced by the peak in the number of counts in the Trans 3 detector’s time spectrum. After µs, they traverse the region of maximum acceptance for the Trans 2 detector, leading to a peak in the Trans 2 time spectrum. Finally, after approximately 5 µs, the muons reach the tip of the target and enter the acceptance region of the Tile 2,5 coincidence resulting in a pronounced increase of counts in the Tile 2,5 time spectrum. The decrease in counts at later times in the Trans detectors’ time spectra is due to muons "flying by" regions with maximal detection efficiency (see Fig. 4). In contrast, the counts in the Tile 2,5 spectra remain steady at later times, indicating that the muons have reached and come to rest at the target tip, where detection efficiency is maximal.
Without transverse (-direction) compression, the majority of the muons would impact the top or bottom target walls before reaching the target tip, which would limit the counts in the Tile 2,5 detectors. Additionally, it would reduce the "fly-by" effect (decrease at late times) in the Trans detectors. The large number of counts in the Tile 2,5 spectrum indicates not only that muons are drifting in -direction but also that rather efficient transverse compression is indeed taking place.
The onset of an effective longitudinal compression is well visible by comparing the Tile time spectra of Fig. 6. Only the spectrum of the Tile 2,5 coincidence, which is sensitive to muons around , shows an increase of counts at late times. The other two pairs, Tile 1,4 and Tile 3,6, which are placed at mm show basically flat time spectra at small count number. This demonstrates that, as muons drift in the -direction, they are efficiently attracted to the central plane of the target at , and do not reach or cross into the acceptance regions of the Tile 1,4 and Tile 3,6 detectors. Together with the substantial increase in counts at late times compared to in Tile 2,5, this provides compelling evidence that efficient longitudinal compression is occurring in the -direction.
@@
Simulations show that such misalignments can considerably influence the observed time spectra. Specifically, the black time spectra in Fig. 5, simulated assuming the initial stop distribution shown in Fig. 7 (Right), display a markedly improved agreement with the measurements. Indeed, by misaligning the stop distribution of the muon beam as shown in Fig. 7 (Right) toward smaller - and -values causes the muons to travel a longer path while moving through a region of higher gas density compared to the properly aligned case. As a result, the time spectra shown in black in Fig. 5 exhibit a shift to later times compared to the red curves of the aligned case. These time spectra were fitted to the measurements using two free parameters for each detector pair (Trans 3, Trans 2, Tile 2,5): the first is a normalization to account for uncertainties in the absolute efficiency of the detector-pairs and the muon stopping probability, the second is a flat (muon-correlated) background to address the effects of muons stopping at the target endcaps and the uncertainty in energy deposition within the small scintillators embedded in the large copper collimators.
Similarly, Fig. 8 presents measured and simulated time spectra for the same three-detector pairs —Trans 3, Trans 2, and Tile 2,5 — but for a 5 mbar target pressure (compared to the 10 mbar of Fig. 5) and for variations of the HV settings. The same misalignment used in Fig. 5 was assumed in the simulations. In this case the two fits in each panel share a common normalization and a common additional muon-uncorrelated background.
@@
Yet, from these measurements, it is possible to roughly estimate the fraction of stopped muons that reach the region of maximal acceptance of the Tile 2,5 detectors. This estimation can be made by comparing the ratio of counts at late times to counts at early times in the Tile 2,5, as obtained from both simulations and measurements. The advantage of this method is that it is largely independent of the muon stopping probability and certain aspects related to detection efficiency. The ratio obtained from simulations closely matches the ratio from the measured time spectra, indicating that the simulations and measurements are consistent.
@@ 6 Conclusion
Although the detection system did neither allow for the precise measurement of the compression efficiency nor the determination of the final spot size, the agreement ultimately achieved between simulation and experiment validates both our experimental technique and the simulations, predicting a 90% compression efficiency (excluding muon decay) within 5 µs. This efficiency allows an estimate of the overall compression efficiency of up to , taking into account the coupling efficiency into the solenoid, the stopping probability in the target, muon decay and losses during extraction and re-acceleration.

### v3

@@ 4 Experimental setup
Figure 4 (Top) shows the placement of the various scintillators around the target as side (Left) and top (Right) views. These scintillators (positron detectors) can be classified into two main categories: the transverse detectors (T1 and T2) used to expose the muons’ motion along the -axis, and the L1-L3 detectors that are sensitive to the muons reaching the target tip exposing at the same time longitudinal () and transverse () compressions. For reference the T1 and T2 detectors have a size of () mm3 while the bottom L1-L3 detectors are () mm3.
Figure 4 (Middle) and (Bottom) depict the detection efficiencies of these detectors, simulated with GEANT4, projected onto the and -plane, respectively. The position-dependence of the detection probability, which results in a relatively good position resolution, was obtained by embedding the scintillators into a massive copper collimator and by organizing the scintillators in telescope-pairs read out in coincidence (top and bottom detectors). Each detector effectively covers distinct zones of the target volume. For instance, the L1-L3 detectors are only sensitive to muon-decays occurring at the tip of the target at their respective -position. For reference Fig. 4 also indicates the position of the muons at time , i.e., the muon stopping distribution before compression takes place. To further increase the position resolution an energy threshold is applied to each positron counter, efficiently removing events from secondary particles. Each scintillator was optically coupled to a wavelength-shifting fiber to transport their optical signals from the cryogenic section of the setup to outside the cryostat. The fibers were then read out by SiPMs [28].
@@ 5 Measured time spectra and comparison to simulation
The time spectra shown in Fig. 5 clearly prove that starting from the muon stopping distribution centered at mm (see Fig. 4) at , the muons gradually drift in the -direction over time. At , the muons are located in the muon stopping region, which falls outside the detection acceptance of the various detectors, resulting in zero or nearly zero counts in all three time spectra. Around µs, the muons pass through the region with the maximum acceptance for the T1 coincidence, as evidenced by the peak in the number of counts in the T1 coincidence time spectrum. After µs, they traverse the region of maximum acceptance for the T2 coincidence, leading to a peak in the T2 time spectrum. Finally, after approximately 5 µs, the muons reach the tip of the target and enter the acceptance region of the L2 coincidence resulting in a pronounced increase of counts in the L2 time spectrum. The decrease in counts for in the T1 time spectrum is due to muons leaving the region of maximal detection efficiency for the T1 coincidence, which is located around mm for (see Fig. 4). A similar decrease is observed in the T2 time spectrum for , indicating that muons pass through and eventually leave the region of maximal detection efficiency for the T2 coincidence, located around mm. In contrast, the lifetime-compensated counts in the L2 spectrum remain constant at late times. Indeed, as shown in Fig. 4 (Right), muons drifting in the -direction are unable to exit the region of highest detection efficiency (for the L2 coincidence) located at mm, as the target ends up at mm.
At late times, all muons must eventually come to rest on one of the target walls, such that the lifetime-compensated time spectrum becomes flat. Without transverse (i.e., -direction) compression, most muons would hit the top or bottom target walls before reaching the tip of the target, resulting in much less L2 coincidence counts and a less pronounced decrease in the T1 and T2 signals. The large number of counts observed in the L2 spectrum therefore demonstrates not only that muons are drifting in the -direction, but also that an efficient transverse compression is indeed taking place.
The onset of an effective longitudinal compression is well visible by comparing the L1-L3 time spectra of Fig. 6. Only the spectrum of the L2 coincidence, which is sensitive to muons around , shows an increase of counts at late times. The other two pairs, L1 and L3, which are placed at mm show basically flat time spectra at small count number. This demonstrates that, as muons drift in the -direction, they are efficiently attracted to the central plane of the target at , and do not reach or cross into the acceptance regions of the L1 and L3 detectors. Together with the substantial increase in counts at late times compared to in the L2, this provides compelling evidence that efficient longitudinal compression is occurring in the -direction as expected from simulations.
@@
Simulations show that such misalignments can considerably influence the observed time spectra. Specifically, the black time spectra in Fig. 5, simulated assuming the initial stop distribution shown in Fig. 7 (Right), display a markedly improved agreement with the measurements. Indeed, by misaligning the stop distribution of the muon beam as shown in Fig. 7 (Right) toward smaller - and -values causes the muons to travel a longer path while moving through a region of higher gas density compared to the properly aligned case. As a result, the time spectra shown in black in Fig. 5 exhibit a shift to later times compared to the red curves of the aligned case. These time spectra were fitted to the measurements using two free parameters for each detector pair (T1, T2, and L2): the first is a normalization to account for uncertainties in the absolute efficiency of the detector-pairs and the muon stopping probability, the second is a flat (muon-correlated) background to address the effects of muons stopping at the target endcaps and the uncertainty in energy deposition within the small scintillators embedded in the large copper collimators.
Similarly, Fig. 8 presents measured and simulated time spectra for the same three-detector pairs —T1, T2, and L2 — but for a 5 mbar target pressure (compared to the 10 mbar of Fig. 5) and for variations of the HV settings. The same misalignment used in Fig. 5 was assumed in the simulations. In this case the two fits in each panel share a common normalization and a common additional muon-uncorrelated background.
@@
Yet, from these measurements, it is possible to roughly estimate the fraction of stopped muons that reach the region of maximal acceptance of the L2 coincidence. This estimation can be made by comparing the ratio of counts at late times to counts at early times in the L2 time spectra, as obtained from both simulations and measurements. The advantage of this method is that it is largely independent of the muon stopping probability and certain aspects related to detection efficiency. The ratio obtained from simulations closely matches the ratio from the measured time spectra, indicating that the simulations and measurements are consistent.
@@ 6 Conclusion
Although the detection system did neither allow for the precise measurement of the compression efficiency nor the determination of the final spot size, the agreement ultimately achieved between simulation and experiment validates both our experimental technique and the simulations, predicting a 90% compression efficiency (excluding muon decay) within 5 µs. This efficiency allows us to estimate the overall performance of the muCool compression scheme. The total efficiency is estimated by considering each sequential stage — starting from muon injection into the solenoid and target, followed by compression, extraction, and final re-acceleration — with the following approximate efficiencies for each step as obtained from GEANT4 simulations: off-axis injection into the solenoid (); transmission through the target entrance window (); muon stopping probability in the active region of the helium gas target (); compression efficiency, excluding muon decay (); survival during the compression time, accounting for muon decay (); extraction efficiency, excluding muon decay (); survival during the time needed for extraction and propagation to the re-acceleration stage (); coupling into the re-acceleration stage (); and re-acceleration to followed by extraction from the magnetic field (). Multiplying these factors yields an overall efficiency of . This estimate assumes operation with the HIMB beam [20, 21] at a momentum of and a target performance as obtained in this study, operated at 7 mbar. One of the dominant uncertainties arises from the extraction from the magnetic field, conservatively estimated at . This process has so far only been studied at a very preliminary level and depends critically on the beam quality at the ferromagnetic plate. A more reliable estimate will require, among other steps, the implementation of dedicated collisional processes in the re-acceleration region in the keV regime, as well as detailed studies of the lenses that must refocus the beam after the position-dependent azimuthal re-acceleration at the ferromagnetic plate [36]. The overall efficiency can be improved by enhancing the target performance, for example by enlarging the active region, optimizing the electric field, refining the gas density distribution, and employing an array of muCool targets. Importantly, the overall efficiency is highly input-beam-dependent: replacing the HIMB beam with the beam boosts it by a factor of five. This boost arises from the smaller momentum bite (enhancing stopping probability), reduced transverse momentum (minimizing reflections in the solenoid), and smaller beam size (improving transmission through the entrance window) of the beam. Efficiencies up to therefore appear within reach for the muCool scheme.